Common Name  Species  Scoville Units
Bell Capsicum annuum  0
Cherry Capsicum annuum  100−500
Pepperoncini Capsicum annuum  100−500
Ancho (Poblano) Capsicum annuum  1000−1500
Jalapeño Capsicum annuum  2500−5000
Yellow Wax (Banana) Capsicum annuum  5000−10000
Serrano Capsicum annuum  10000−23000
Cayenne Capsicum annuum  30000−50000
Tobasco Capsicum frutescens  30000−50000
Habañero Capsicum chinense  100000−300000

Table 1: Scoville Units of popular members of the Capsicum genus